Andre traveled from his home to the office at an average rate of 25 kph. By traveling 5 kph faster, he took 30 min less to return home. How far is his office from his home?
Answer by checkley77(12844)

D=25T
D=30(T-.5))
25T=30(T-.5)
25T=30T-15
25T-30T=-15
-5T=-15
T=-15/-5
T=3 HOURS FOR THE 25MPH. TRIP.
25*3=75 MILES FROM HOME TO OFFICE
30(3-.5)=30*2.5=75 DITTO.